# Towards Trapped-Ion Thermometry Using Cavity-Based EIT

### v1

@@ Abstract
We present a technique for measuring ion temperature using cavity-based electromagnetically induced transparency (EIT) applicable for cavity-qed systems in the strong coupling regime. This method enables efficient extraction of the ion’s phonon occupation number following sub-Doppler cooling close to the motional ground state. The proposed method relies on monitoring the cavity probe transmission while scanning the probe laser frequency once cavity EIT is established using the control beam, significantly simplifying the measurement procedure. We theoretically establish a model that demonstrates influence of thermal state of the trapped ion vis a vis the EIT linewidth measured. We show how the cavity EIT transmission may be used as a thermometry tool to deduce the ion temperature as well as the motional state for an ion in the sub-Doppler cooling regime. The current method can only be used for operation in the resolved-sideband regime, where individual motional states can be selectively addressed for all relevant transitions either by selecting appropriate energy levels for the three-level system or by employing strong confinement with high secular frequencies ($\sim 10MHz$).
@@ I Introduction
The paper is organized as follows: In section II, we present the theoretical model and Hamiltonian including the treatment of ion motion and dissipative processes. In section III, numerical results for linewidth extraction and cavity transmission are shown. The section IV examines how phonons affect the cavity-EIT spectrum. The analysis of ion temperature and the role of thermalization dynamics are discussed in section V. Finally, a summary of the key results and concluding remarks are given in section section VII.
@@ V Determination of ion temperature and average phonon number
VI Roadmap towards experimental realization
We present here an outline of experimentally implementing the trapped-ion thermometry method presented in this paper. As the thermal state detection method proposed is cavity EIT-based, one needs to engineer a trapped-ion cavity-QED system in the strong-coupling regime ($g^{2}\gg\kappa\gamma$). The cavity configuration may be chosen to be concentric or confocal as per the requirements of the physics objectives. We recommend using a high finesse (>50000>50000) cavity, such that the coupling factor ($g$) is high enough to justify operation in the strong coupling regime ($g^{2}\gg\kappa\gamma$).
@@ VI Roadmap towards experimental realization
One may also consider using a dipole allowed transition (where typical linewidths are in the range of $\approx 5-20MHz$) for the trapped ion strongly coupled to the cavity field, one should consider operating the ion trap in a very tight confinement (very deep trap potential) of the ion with secular frequency greater transition linewidth ($\Gamma$). Such a mode of ion trap operation has been experimentally validated [35], where an 9Be+ ion was trapped with a very high RF voltage (600 V) and frequency (240 MHz), reaching secular frequencies in the range of $\sim 50~MHz$ and the sideband transitions were addressed in a dipole allowed transition .
VII Conclusion
@@ Acknowledgments
A two-dimensional map of the cavity-EIT linewidth represented as the normalized full width at half maxima (FWHM) ratio as a function of the normalized control field Rabi frequency $\Omega_{c}$ and the normalized atom-cavity coupling strength $g$ for various mean thermal phonon numbers $n_{th}$ is shown in Figure 5 For a weakly excited motional state ($n_{th}=0.5$) (see 5(a)), the linewidth ratio remains close to unity over a large region of the parameter space. This suggests that thermal motion only slightly alters EIT coherence. As the value of thermal occupation increases $n_{th}=1$ (see 5(b)), a clear reduction of the FWHM ratio occurs, particularly in the regime of strong ion-cavity coupling and weak control fields. This reduction signals the onset of thermally induced decoherence. A higher phonon number ($n_{th}=5,10$) (see 5(c), 5(d)) shows a more noticeable trend in which the parameter space is dominated by extended regions of reduced linewidth ratio. In this regime, thermal motion significantly perturbs the ion-cavity interaction, leading to enhanced dephasing and a degradation of the EIT interference.

### v2

@@ Abstract
We present a technique for measuring ion temperature using cavity-based electromagnetically induced transparency (EIT) applicable for cavity-qed systems in the strong coupling regime. This method enables efficient extraction of the ion’s phonon occupation number following sub-Doppler cooling close to the motional ground state. The proposed method relies on monitoring the cavity probe transmission while scanning the probe laser frequency once cavity EIT is established using the control beam, significantly simplifying the measurement procedure. We theoretically establish a model that demonstrates influence of thermal state of the trapped ion vis-a-vis the EIT linewidth measured. We show how the cavity EIT transmission may be used as a thermometry tool to deduce the ion temperature as well as the motional state for an ion in the sub-Doppler cooling regime even for systems that are in the weak coupling regime. The current method can only be used for operation in the resolved-sideband regime, where individual motional states can be selectively addressed for all relevant transitions either by selecting appropriate energy levels for the three-level system or by employing strong confinement with high secular frequencies ($\sim 10MHz$).
@@ I Introduction
The paper is organized as follows: In section II, we present the theoretical model and Hamiltonian including the treatment of ion motion and dissipative processes. In section III, numerical results for linewidth extraction and cavity transmission are shown. The section IV examines how phonons affect the cavity-EIT spectrum. The analysis of ion temperature and the role of thermalization dynamics are discussed in section V. Finally, a summary of the key results and concluding remarks are given in section section VIII.
@@ V Determination of ion temperature and average phonon number
VI Thermometry of a multi-ion system using a low finesse cavity
The discussion we have upto this point considers a single ion confined in a high-finesse cavity and operating in a strong coupling regime. However, operating in a strong-coupling regime is experimentally challenging with only a single ion, requiring either a near concentric-cavity regime [35] or fiber cavities where the polished surfaces are placed very close to each other [36]. Interestingly, both of these configurations possess very small mode volumes increasing the single atom cooperativity factor in a significant manner. Both are experimentally very challenging, with substantial efforts involved in making and aligning cavity mirrors with very high reflectivity.
Figure 5: Cavity-EIT linewidth variation with the number of ions. The simulation parameters are shown in the inset.
Even in a weakly coupled cavity regime ($g\ll\kappa,\gamma$), it is possible to apply the proposed cavity EIT-based ion-thermometry method discussed earlier, albeit using more than one ion coupled to the cavity mode instead of a single one. The method works as one may achieve the collective strong coupling regime with a larger number of ions coupled to the cavity mode, thanks to the scaling of the single-atom cavity coupling factor ($g_{eff}\propto\sqrt{N}$) $g$ with ion number, $N$ [37]. Here g is the single-ion cavity coupling factor. The dependence of ion number on cavity-EIT linewidth is depicted in Figure 5. As the ion number increases, the cavity-EIT linewidth decreases exponentially, making the EIT linewidth even narrower than the single-atom case in the strong coupling regime. It suggests that using a sufficient number of atoms/ions collectively coupled to the cavity, one may be able to realize a very narrow cavity-EIT linewidth, which may be interesting for different other applications as well, such as using them as reference transitions in atomic clocks, laser frequency stabilization, etc.
Figure 6: Temperature vs cavity-EIT linewidth for multi-atoms. The EIT linewidth is scaled with respect to the empty cavity width. The typical single-ion cavity EIT linewidth for one temperature is shown in the inset with the set of different parameters. Except temperature the same parameter has been chosen for the simulations performed in this plot.
In the multi-atom/ion cavity system, the ion-cloud temperature may be estimated in a similar manner as outlined in the single ion cavity-EIT based thermometry scheme as discussed in section V. We have shown the variation in temperature as a function of cavity-EIT linewidth in fraction of empty cavity linewidth for different number of ions in Figure 6. The variation of the cavity EIT linewidth with temperature indicates that the linewidth is more sensitive to the temperature as one considers a higher number of atoms/ions coupled to the cavity mode. Thus, the proposed thermometry method is particularly well suited for implementation in low ion-cavity coupling strength cavity qed systems with multiple ions coupled to the cavity mode.
The current methodology discussed for the cavity EIT-based thermometry for a multi-ion cavity qed system has been developed with one simplifying assumption incorporated in our model. We assume that all the atoms/ions coupled to cavity mode possess the same thermal distribution or thermal occupancy. Modeling the system whereby different atoms/ions have a distinct thermal occupation is potentially challenging and computationally expensive. Additionally, it does not change the physics of the phenomena being modeled. Taking all ions to possess the same motional state is a thus a realistic assumption when all the ions are cooled to sufficiently low temperature ($\approx 10-50~\mu K$) and all of them share the same effective coupling to the cavity mode.
VII Roadmap towards experimental realization
We present here an outline of experimentally implementing the trapped-ion thermometry method presented in this paper. As the thermal state detection method proposed is cavity EIT-based, one needs to engineer a trapped-ion cavity-QED system in the strong-coupling regime ($g^{2}\gg\kappa\gamma$). This may be achieved in two ways. The first approach would be to engineer a cavity configuration in the near-concentric or near-confocal regime as per the requirements of the physics objectives. We recommend using a high finesse (>50000>50000) cavity, such that the coupling factor ($g$) is high enough to justify operation in the strong coupling regime ($g^{2}\gg\kappa\gamma$) for a single atom/ion coupled to the cavity mode. The second approach focuses on a relatively relaxed regime of multi-ion based strong coupling regime of operation for a low finesse cavity.
@@ VII Roadmap towards experimental realization
One may also consider using a dipole allowed transition (where typical linewidths are in the range of $\approx 5-20MHz$) for the trapped ion strongly coupled to the cavity field, one should consider operating the ion trap in a very tight confinement (very deep trap potential) of the ion with secular frequency greater transition linewidth ($\Gamma$). Such a mode of ion trap operation has been experimentally validated [38], where an 9Be+ ion was trapped with a very high RF voltage (600 V) and frequency (240 MHz), reaching secular frequencies in the range of $\sim 50~MHz$ and the sideband transitions were addressed in a dipole-allowed transition.
If one considers applying the proposed thermometry method with multiple ions in low to medium finesse (finesse values in the range 3000150003000-15000) cavities having a relatively large mode waist ($50-100~\mu m$), about 10 ions coupled to the cavity mode will be sufficient to reach the strong coupling regime, assuming a single ion cavity coupling factor of $0.5$ MHz, cavity decay rate as well as the excited state decay rate about $1$ MHz.
VIII Conclusion
@@ Acknowledgments
A two-dimensional map of the cavity-EIT linewidth represented as the normalized full width at half maxima (FWHM) ratio as a function of the normalized control field Rabi frequency $\Omega_{c}$ and the normalized atom-cavity coupling strength $g$ for various mean thermal phonon numbers $n_{th}$ is shown in Figure 7 For a weakly excited motional state ($n_{th}=0.5$) (see 7(a)), the linewidth ratio remains close to unity over a large region of the parameter space. This suggests that thermal motion only slightly alters EIT coherence. As the value of thermal occupation increases $n_{th}=1$ (see 7(b)), a clear reduction of the FWHM ratio occurs, particularly in the regime of strong ion-cavity coupling and weak control fields. This reduction signals the onset of thermally induced decoherence. A higher phonon number ($n_{th}=5,10$) (see 7(c), 7(d)) shows a more noticeable trend in which the parameter space is dominated by extended regions of reduced linewidth ratio. In this regime, thermal motion significantly perturbs the ion-cavity interaction, leading to enhanced dephasing and a degradation of the EIT interference.